OpenStudy (anonymous):

Will give a medal ----------------- Which description best describes the solution to the following system of equations? y = −one halfx + 9 y = x + 7 Lines y = −one halfx + 9 and y = x + 7 intersect the x-axis. Lines y = −one halfx + 9 and y = x + 7 intersect the y-axis. Line y = −one halfx + 9 intersects the origin. Line y = −one halfx + 9 intersects line y = x + 7.

OpenStudy (anonymous):

@Decarr432

OpenStudy (anonymous):

@mazmaz1

OpenStudy (ijlal):

i would say that the 4th option is correct that the line y=(-1/x) +9 intersects the line y=x+7 the red graph is of y=(-1/x)+9 and the other is of y=x+7

OpenStudy (anonymous):

Could you help me with another one?

OpenStudy (ijlal):

the first 2nd and third option i found wrong is because i think you mean by saying " Lines y = −one halfx + 9 and y = x + 7 intersect the x-axis. Lines y = −one halfx + 9 and y = x + 7 intersect the y-axis" that they both intersect at the x-y axes but they dont if you see the graph and the third is wrong because y=-(1/x)+9 has 2 ends one going up and one going down so they both dont intersect at origin. So that is why the 4th option is right :)

OpenStudy (ijlal):

@blackinevitability Sure ask away :)

OpenStudy (anonymous):

A set of equations is given below: Equation C: y = 6x + 9 Equation D: y = 6x + 2 Which of the following options is true about the solution to the given set of equations? One solution No solution Two solutions Infinite solutions

OpenStudy (anonymous):

I think its infinite but im not sure

OpenStudy (ijlal):

as both the lines have same gradient it means they are parallel lines remember the property m1=m2 when the lines are parallel so as you see by the graph since the lines are parallel they wont intersect each other at any point which means they don't have any solution.

OpenStudy (anonymous):

OOH Okay, thank you for your help!

OpenStudy (ijlal):

Solution means when the graph intersect at each other and that point at which they intersect is the solution so these lines don't intersect hence they don't have any solution :) so option 2 is right :)

OpenStudy (ijlal):

Your most welcome ;)